OpenStudy (anonymous):

Just a quick question to refresh my memory... Is sin^2(t) the same as [sin(t)]^2 or would I have to switch it to [1/csc(t)]^2 to make it workable in an equation?

OpenStudy (anonymous):

Clarification on the formatting: \[\sin^2\theta\] vs \[(\sin \theta)^2\]

OpenStudy (anonymous):

[sin(t)]^2 1-cos^2(t) (1/2)(1-cos(2t))